# Nonparametric estimation of the jump-size distribution for an M/G/1 storage system with Poisson sampling

### v3

@@ Abstract
This work presents a non-parametric estimator for the cumulative distribution function (CDF) of the jump-size distribution for a storage system with compound Poisson input. The workload process is observed according to an independent Poisson sampling process. The nonparametric estimator is constructed by first estimating the characteristic function (CF) and then applying an inversion formula. The convergence rate of the CF estimator at $s$ is shown to be of the order of $s^{2}/n$, where $n$ is the sample size. This convergence rate is leveraged to explore the bias-variance tradeoff of the inversion estimator. It is demonstrated that within a certain class of continuous distributions, the risk, in terms of MSE, is uniformly bounded by $Cn^{-\frac{\eta}{1+\eta}}$, where $C$ is a positive constant and the parameter $\eta>0$ depends on the smoothness of the underlying class of distributions. A heuristic method is further developed to address the case of an unknown rate of the compound Poisson input process.
@@ 1 Introduction
We consider a system where iid jobs with a continuous distribution $G$ arrive according to a Poisson process with rate $\lambda$ and are processed at a unit rate. This model is known as the canonical M/G/1 queue, although it also often used to model other processes, e.g., risk processes and communication networks. The job-size distribution is crucial in determining the system’s performance. For instance, the mean sojourn time is determined by the first two moments of $G$ (see [2]). However, in many practical applications, the job-size distribution is unknown and needs to be estimated from aggregate data. This work presents a nonparametric estimator for the jump size CDF $G$ from discrete observations of the workload, i.e., a reflected version of the compound Poisson input process with an additional drift term. Worst-case convergence rate guarantees are established for the risk associated with the estimator by combining asymptotic methods drawing inspiration from those applied in statistical deconvolution problems (e.g., [18]) together with the ergodic properties of the workload process (e.g., [36]).
@@ 1.1 Model and objective
Jobs arrive at a single-server queue according to a Poisson process $(N(t))_{t\geq 0}$ with rate $\lambda$. Each job has an independent size with a common continuous distribution $\mathrm{P}(B\leq x)=G(x)$ and CF characteristic function $\gamma(s)=\mathrm{E}[\mathrm{e}^{isB}]$. The input process is therefore a compound Poisson process, denoted by $J(t)=\sum_{i=1}^{N(t)}B_{i}$. Let $X(t)=J(t)-t$ denote the net-input process. The workload process $(V(t))_{t\geq 0}$, representing the virtual waiting time of a potential arrival at time $t\geq 0$,is the net input process reflected at zero: $V(t)=X(t)+\max\{V(0),L(t)\}$, where $L(t):=-\inf_{0\leq s\leq t}X(s)$ (see [2]). The characteristic exponent of the net input process is
@@
Figure 2: The estimators are based on a simulation of $n=10^{4}$ workload observations with sampling rate $\xi=1$. The arrival rate is $\lambda=1$ and the true CDF $G$ (black solid line) is a mixture of gamma distributions $\bm{\alpha}=(2,6,1)$, $\bm{\beta}=(3.5,90,0.09)$ and $\bm{p}=(0.6,0.35,0.05)$. The estimated functions $\widehat{G}_{n}^{(h)}$ are plotted for $h\in\{0.5,1,2,5\}$.
@@
Note that here we used the PASTA property; the limit at Poisson sampling moments coincides with the time-average limit (see [2, Ch.ZVI]). If $\left|\mathrm{E}[\mathrm{e}^{isV}]\right|>\epsilon$, then the estimator is only truncated a finite number of times, almost surely, and the asymptotic performance of the truncated estimator $\widehat{\gamma}^{\epsilon}_{n}(\cdot)$ coincides with that of the original estimator $\widehat{\gamma}_{n}(\cdot)$. In Section 2 we further discuss how the choice of $\epsilon$ is related to specific assumptions on the class of possible job-size distributions.
@@ 1.2 Main contributions
    where $h_{n}=n^{1/(2(1+\eta)}$. This further provides a practical way to select the truncation parameter $h$ for a given sample size. The class of functions $\mathcal{G}$ includes the ordinary-smooth and super-smooth distributions in the terminology of the statistical deconvolution literature (see [18]). Examples of such distributions are Gamma, log-normal, truncated normal, pareto and their mixtures. Note that this result provides an upper bound on the risk, but does not guarantee rate-optimality, in the sense of [35]. Establishing a tight lower bound on the risk seems to be a very challenging problem due to the intractability of the joint distribution of the workload observations. Nevertheless, this work is the first to present any type of guarantee on the rate for nonparametric estimation of $G$.
@@
  • The theoretical results are supplemented with simulation analysis and a discussion of practical implementation challenges. In particular, we examine the sensitivity of the procedure to the truncation parameter $h$. While in the asymptotic regime, $h_{n}$ must tend to infinity for the bias to vanish, we observe that for finite samples, moderate values of $h$ outperform larger ones. This occurs because the variance of the CF estimator is $O(s^{2}/n$, so increasing $h$ incorporates noisier terms into the integral of the inversion formula.
@@ 1.3 Related literature
The first works on this topic naturally focused on parametric methods, some of which are surveyed in [11]. For example, [8] considered a single-server queue that is observed for some fixed duration of time. This results in a random sample of arrival and service times following some parametric distributions, for which an MLE is computed. The large sample asymptotic performance of the estimator is further studied. In [7] an MLE is derived for a G/G/1 queue with waiting time observations by assuming the increment in Lindley’s recursive formula follows a parametric distribution. For estimators based on waiting time (or queue length) data, asymptotic analysis of must take into account the covariance structure of the process (see [32]), and this will also play a key role in the analysis presented here. Another key feature in statistical inference for queueing systems is that it typically involves indirect estimation. For example, in [8] waiting times observations are treated as a censored version of the increments in Lindley’s formula, i.e., censoring occurs when the waiting time is zero.
@@
The challenges of indirect estimation related to probing data are considered in detail in [6], and estimators for specific models are discussed. A nonparametric estimator for $G$ based on discrete workload observations is provided in [23]. Specifically, Pollaczeck-Khintchine’s formula is inverted to obtain an estimator of the residual job-size distribution from the empirical workload distribution. However, estimating $G$ from its residual counterpart requires a heuristic step that is shown to be very noisy for systems with very high or very low load. Nonparametric inference of the job-size distribution is mostly carried out for infinite server queues and networks (e.g., [10], [33] [19] and [20]).
@@
The problem of decompounding a compound Poisson sum was studied in [12]. In this setting the goal is to infer $G$ by observing realizations of $\sum_{i=1}^{N(t)}B_{i}$. Inversion techniques have been applied for this problem are explored in [17] and a Bayesian approach is presented in [21]. Our setting is distinct from the decompounding problem in that we do not observe the compound Poisson process, but rather a process that includes a drift term and is reflected at zero. An inversion estimator for a similar process without reflection, namely a finite intensity subordinator (compound Poisson process with drift), that is observed at discrete time points is given in [22]. A rate optimal density estimator is derived in [16] for the jumps of a discretely observed compound Poisson process (without reflection). Estimation of the characteristic exponent of general Lévy processes from discrete (low frequency) observations is further studied in [13], [28] and [9]. The methods applied here are also related to those of nonparametric noise deconvolution problems. In particular, the rate of convergence of the estimator is related to the smoothness class of the job-size distribution (see [18] and [14]).
@@ 2 Main results
$\mathcal{R}_{\mathcal{G}}(\widehat{G};x)=\sup_{G\in\mathcal{G}}R(\widehat{G},G;x)$
@@ 2.1 On the risk convergence rate
If $\eta$ is very high, i.e. the distribution is ‘very smooth’, then the convergence rate is almost parametric ($1/n$). In general, the value $1/\eta$ can be thought of as a penalty to the parametric rate that is higher as the distribution is ‘less smooth’. It is interesting to point out that very similar convergence rates, i.e., polynomial that are close to parametric when $\eta$ is low, are shown to be rate-optimal in the setting of kernel density estimation with additive observation errors that satisfy condition (13) (e.g., [18, Thm. 2]). Of course, in practice $\eta$ is not known and setting a specific value for $h_{n}=n^{1/(2(1+\eta))}$ represents some prior belief on the smoothness of the possible distributions. In particular, choosing a low value of $\eta$ can be considered conservative. In the next examples we illustrate this interplay for specific distributions. The examples highlight the fact that as the underlying smoothness of $G$, indexed by the parameter $\eta$, increases, the estimation accuracy improves and a lower truncation parameter $h$ is required.
@@
(Continued) Mixture of Gamma distributions: In Section 1.1 the estimator was applied with different choices of $h$ to a mixture of Gamma distributions with $\bm{\alpha}=(2,6,1)$, $\bm{\beta}=(3.5,90,0.09)$ and $\bm{p}=(0.6,0.35,0.05)$. Theorem 2 prescribes selecting the minimal shape parameter $\eta=\min\bm{\alpha}=1$, yielding $\zeta=1/4$ and $h_{n}\approx 10$. In Figure 2 the estimated functions are plotted for a sample size of $n=10^{4}$. It was observed in the example that the best fit appears to be in the range of $[5,10]$. Keep in mind that the rates in Theorem 2 are valid up to scaling by constants so $h_{n}\approx 10$ is certainly in line with the empirical observations for this example. The MSE upper bound (15) for $n=10^{4}$ is of of the order $n^{-1/2}=0.01$ for $h_{n}=10$. Of course, these bounds are scaled by constants which may still be significant for this sample size. $\diamond$
@@
Log-normal distribution: Suppose now that the job-sizes follow a log-Normal distribution with parameters $(\mu,\sigma)$; $G(x)=\Phi((\log x-\mu)/\sigma)$. The characteristic function does not admit a closed form expression but via approximations it can be shown that it decays in absolute value at a polynomial rate with $\eta>2$ (see [3]). For the case of $\mu=0.2$ and $\sigma=0.5$ the estimator was applied for different $h$ for a sample of $n=10^{4}$ workload observations. Hence, by Theorem 2 we expect that $h_{n}=n^{1/6}\approx 4.6$ will yield a good fit. The results are illustrated in Figure 3. Indeed, the best fit is achieved with $h=4$, while increasing the truncation threshold yields a noisier estimator. $\diamond$
@@ 3.2 Convergence of the covariance terms
First note that $\mathrm{P}(V_{0}<\infty)=1$ when $V_{0}$ follows the stationary distribution. For an M/G/1 process, it is known that if $\lambda\mathrm{E}[B]<1$ and $\mathrm{E}[B^{k+1}]<\infty$, then the workload process at arrival moments is ergodic with a convergence rate of $n^{k}$ (see [36]). We will show that this result holds for the workload process sampled according to a Poisson process. To this end we first construct a discrete-time version of the workload process at arrival and sampling times. Due to the memoryless property, at any state of the system the time until the next event of a departure or an observation is an exponential random variable with rate $\lambda+\xi$. Let $(\tau_{k})_{k\geq 0}$ denote an iid sequence of such inter-event times. Let $I_{k}$ denote the indicator of event $k$ being a an arrival event, then $I_{k}$ is a Bernoulli random variable with probability $\lambda/(\lambda+\xi)$, and it is also independent of $\tau_{k}$. Let $\tilde{V}_{k}:=\lim_{t\downarrow 0}V(\sum_{j=1}^{k}\tau_{j}-t)$ denote the workload just before the $k$’th event. The standard Lindley recursion yields
@@
for any bounded function $g$ and initial state $V_{0}\in[0,\infty)$. Let $S(k)$ denote the index of the $k$’th sampling event in the sequence of events driving $(\tilde{V})_{n\geq 1}$, and note that $S(k)-S(k-1)$ follows a geometric distribution with parameter $\xi/(\lambda+\xi)$. By the SLLN, $S(k)/k\>\to_{\rm as}\>\xi/(\lambda+\xi)$ as $k\to\infty$. Clearly, $S(k)\geq k$, and in particular $S(k)\>\to_{\rm as}\>\infty$ as $k\to\infty$. Thus, (38) implies that
@@
Applying L’Hôpital’s rule rule three times to $\kappa_{1}(s_{1},s_{2};G)$ to establish that the limit as $s_{2}\to 0$ is bounded under the following condition,
@@
$\kappa_{2}(s;G)$ $=\sum_{k=1}^{\infty}\left(\mathrm{E}[\mathrm{e}^{isV_{0}}\mathbf{1}\{V_{k}=0\}-\mathrm{E}[\mathrm{e}^{isV}]\mathrm{P}(V=0)\right)=\mathrm{E}[\mathrm{e}^{isV}]\mathrm{E}\left[\sum_{k=1}^{\infty}\left(\mathrm{P}(V_{k}=0|V_{0})-\mathrm{P}(V=0)\right)\right]$

### v4

# Nonparametric estimation of the job-size distribution for an M/G/1 queue with Poisson sampling [^1] To appear in Queueing Systems: Theory and Applications

@@ Abstract
This work presents a non-parametric estimator for the cumulative distribution function (CDF) of the job-size distribution for a queue with compound Poisson input. The workload process is observed according to an independent Poisson sampling process. The nonparametric estimator is constructed by first estimating the characteristic function (CF) and then applying an inversion formula. The convergence rate of the CF estimator at $s$ is shown to be of the order of $s^{2}/n$, where $n$ is the sample size. This convergence rate is leveraged to explore the bias-variance tradeoff of the inversion estimator. It is demonstrated that within a certain class of continuous distributions, the risk, in terms of MSE, is uniformly bounded by $Cn^{-\frac{\eta}{1+\eta}}$, where $C$ is a positive constant and the parameter $\eta>0$ depends on the smoothness of the underlying class of distributions. A heuristic method is further developed to address the case of an unknown rate of the compound Poisson input process.
@@ 1 Introduction
We consider a system where iid jobs with a continuous distribution $G$ arrive according to a Poisson process with rate $\lambda$ and are processed at a unit rate. This model is known as the canonical M/G/1 queue, although it also often used to model other processes, e.g., risk processes and communication networks. The job-size distribution is crucial in determining the system’s performance. For instance, the mean sojourn time is determined by the first two moments of $G$ (see [2]). However, in many practical applications, the job-size distribution is unknown and needs to be estimated from aggregate data. This work presents a nonparametric estimator for the jump size cumulative distribution function (CDF) $G$ from discrete observations of the workload, i.e., a reflected version of the compound Poisson input process with an additional drift term. Worst-case convergence rate guarantees are established for the risk associated with the estimator by combining asymptotic methods drawing inspiration from those applied in statistical deconvolution problems (e.g., [18]) together with the ergodic properties of the workload process (e.g., [36]).
@@ 1.1 Model and objective
Jobs arrive at a single-server queue according to a Poisson process $(N(t))_{t\geq 0}$ with rate $\lambda$. Each job has an independent size with a common continuous distribution $\mathrm{P}(B\leq x)=G(x)$ and CF characteristic function $\gamma(s)=\mathrm{E}[\mathrm{e}^{isB}]$. The input process is therefore a compound Poisson process, denoted by $J(t)=\sum_{i=1}^{N(t)}B_{i}$. Let X(t)=J(t)tX(t)=J(t)-t denote the net-input process. The workload process $(V(t))_{t\geq 0}$, representing the virtual waiting time of a potential arrival at time $t\geq 0$,is the net input process reflected at zero: $V(t)=X(t)+\max\{V(0),L(t)\}$, where $L(t):=-\inf_{0\leq s\leq t}X(s)$ (see [2]). The characteristic exponent of the net input process is
@@
Figure 2: The estimators are based on a simulation of $n=10^{4}$ workload observations with sampling rate $\xi=1$. The arrival rate is $\lambda=1$ and the true CDF $G$ (black solid line) is a mixture of gamma distributions $\bm{\alpha}=(2,6,1)$, $\bm{\beta}=(3.5,90,0.09)$ and $\bm{p}=(0.6,0.35,0.05)$. The estimated functions $\widehat{G}_{n}^{(h)}$ are plotted for $h\in\{1,2,5,10\}$.
@@
Note that here we used the PASTA property; the limit at Poisson sampling moments coincides with the time-average limit (see [2, Ch.ZVI]). If $\left|\mathrm{E}[\mathrm{e}^{isV}]\right|>\epsilon$, then the estimator is only truncated a finite number of times, almost surely, and the asymptotic performance of the truncated estimator $\widehat{\gamma}^{\epsilon}_{n}(\cdot)$ coincides with that of the original estimator $\widehat{\gamma}_{n}(\cdot)$. The choice of $(1-\epsilon)(1+i)$ ensures that the truncated CF estimator lies in $(0,1)$, although, in principle other constants may be used without affecting the asymptotic analysis. In Section 2 we further discuss how the choice of $\epsilon$ is related to specific assumptions on the class of possible job-size distributions.
@@ 1.2 Main contributions
    where $h_{n}=n^{1/(2(1+\eta))}$. This further provides a practical way to select the truncation parameter $h$ for a given sample size. The class of functions $\mathcal{G}$ includes the ordinary-smooth and super-smooth distributions in the terminology of the statistical deconvolution literature (see [18]). Examples of such distributions are Gamma, log-normal, truncated normal, pareto and their mixtures. Note that this result provides an upper bound on the risk, but does not guarantee rate-optimality, in the sense of [35]. Establishing a tight lower bound on the risk seems to be a very challenging problem due to the intractability of the joint distribution of the workload observations. Nevertheless, this work is the first to present any type of guarantee on the rate for nonparametric estimation of $G$.
@@
  • The theoretical results are supplemented with simulation analysis and a discussion of practical implementation challenges. In particular, we examine the sensitivity of the procedure to the truncation parameter $h$. While in the asymptotic regime, $h_{n}$ must tend to infinity for the bias to vanish, we observe that for finite samples, moderate values of $h$ outperform larger ones. This occurs because the variance of the CF estimator is $O(s^{2}/n)$, so increasing $h$ incorporates noisier terms into the integral of the inversion formula. An interesting observation in these experiments is that the estimation accuracy is better when $\rho$ is lower, but the optimal choice of truncation parameter remains more or less the same. The same is observed for a slower sampling rate, which also reduces the variance of estimation errors because the correlation between workload observations is weaker.
@@ 1.3 Related literature
The first works on this topic naturally focused on parametric methods, some of which are surveyed in [11]. For example, [8] considered a single-server queue that is observed for some fixed duration of time. This results in a random sample of arrival and service times following some parametric distributions, for which an MLE is computed. The large sample asymptotic performance of the estimator is further studied. In [7] an MLE is derived for a G/G/1 queue with waiting time observations by assuming the increment in Lindley’s recursive formula follows a parametric distribution. For estimators based on waiting time (or queue length) data, asymptotic analysis of the estimation error must take into account the covariance structure of the process (see [32]), and this will also play a key role in the analysis presented here. Another key feature in statistical inference for queueing systems is that it typically involves indirect estimation. For example, in [8] waiting times observations are treated as a censored version of the increments in Lindley’s formula, i.e., censoring occurs when the waiting time is zero.
@@
The challenges of indirect estimation related to probing data are considered in detail in [6], and estimators for specific models are discussed. A nonparametric estimator for $G$ based on discrete workload observations is provided in [23]. Specifically, Pollaczek–Khinchine’s formula is inverted to obtain an estimator of the residual job-size distribution from the empirical workload distribution. However, estimating $G$ from its residual counterpart requires a heuristic step that is shown to be very noisy for systems with very high or very low load. Nonparametric inference of the job-size distribution is mostly carried out for infinite server queues and networks (e.g., [10], [33] [19] and [20]).
@@
The problem of decompounding a compound Poisson sum was studied in [12]. In this setting the goal is to infer $G$ by observing realizations of $\sum_{i=1}^{N(t)}B_{i}$. Inversion techniques that have been applied for this problem are explored in [17] and a Bayesian approach is presented in [21]. Our setting is distinct from the decompounding problem in that we do not observe the compound Poisson process, but rather a process that includes a drift term and is reflected at zero. An inversion estimator for a similar process without reflection, namely a finite intensity subordinator (compound Poisson process with drift), that is observed at discrete time points is given in [22]. A rate optimal density estimator is derived in [16] for the jumps of a discretely observed compound Poisson process (without reflection). Estimation of the characteristic exponent of general Lévy processes from discrete (low frequency) observations is further studied in [13], [28] and [9]. The methods applied here are also related to those of nonparametric noise deconvolution problems. In particular, the rate of convergence of the estimator is related to the smoothness class of the job-size distribution (see [18] and [14]).
@@ 2 Main results
$\mathcal{R}_{\mathcal{G}}(\widehat{G};x)=\sup_{G\in\mathcal{G}}R(\widehat{G},G;x)\$
@@ 2.1 On the risk convergence rate
If $\eta$ is very high, i.e. the distribution is ‘very smooth’, then the convergence rate is almost parametric ($1/n$). In general, the value $1/\eta$ can be thought of as a penalty to the parametric rate that is higher as the distribution is ‘less smooth’. It is interesting to point out that very similar convergence rates, i.e., polynomial that are close to parametric when $\eta$ is low, are shown to be rate-optimal in the setting of kernel density estimation with additive observation errors that satisfy condition (13) (e.g., [18, Thm. 2]). Of course, in practice $\eta$ is not known and setting a specific value for $h_{n}=n^{1/(2(1+\eta))}$ represents some prior belief on the smoothness of the possible distributions. In particular, choosing a low value of $\eta$ can be considered conservative. In the next examples we illustrate this interplay for specific distributions. The examples highlight the fact that as the underlying smoothness of $G$, indexed by the parameter $\eta$, increases, the estimation accuracy improves and a lower truncation parameter $h$ is required. The following simulation experiments suggest that selecting a truncation parameter that yields a smoother estimated CDF is a reasonable heuristic for choosing $h$ when $\eta$ is unknown.
@@
(Continued) Mixture of Gamma distributions: In Section 1.1 the estimator was applied with different choices of $h$ to a mixture of Gamma distributions with $\bm{\alpha}=(2,6,1)$, $\bm{\beta}=(3.5,90,0.09)$ and $\bm{p}=(0.6,0.35,0.05)$. Theorem 2 prescribes selecting the minimal shape parameter $\eta=\min\bm{\alpha}=1$, yielding $\zeta=1/4$ and $h_{n}\approx 10$. In Figure 2 the estimated functions are plotted for a sample size of $n=10^{4}$. It was observed in the example that the best fit appears to be in the range of $[5,10]$. Keep in mind that the rates in Theorem 2 are valid up to scaling by constants so $h_{n}\approx 10$ is certainly in line with the empirical observations for this example. The MSE upper bound (15) for $n=10^{4}$ is of of the order $n^{-1/2}=0.01$ for $h_{n}=10$. Of course, these bounds are scaled by constants which may still be significant for this sample size.
In practice, the underlying distribution and $\eta$ itself are unknown, so the above is not a piratical recipe for choosing an appropriate $h$. A useful observation is that that the estimated function with $h=10$ displays much more moderate fluctuations compared to all other choices, thus suggesting that this is a reasonable heuristic criteria for the choice of $h$. The same behavior is observed in the following two examples, further strengthening the usefulness of this heuristic. $\diamond$
@@
Log-normal distribution: Suppose now that the job-sizes follow a log-Normal distribution with parameters $(\mu,\sigma)$; $G(x)=\Phi((\log x-\mu)/\sigma)$. The characteristic function does not admit a closed form expression but via approximations it can be shown that it decays in absolute value at a polynomial rate with $\eta>2$ (see [3]). For the case of $\lambda=0.6$ $\mu=0.2$, and $\sigma=0.5$ the estimator was applied for different $h$ for a sample of $n=10^{4}$ workload observations. Hence, by Theorem 2 we expect that $h_{n}=n^{1/6}\approx 4.6$ will yield a good fit. The results are illustrated in Figure 3. Indeed, the best fit is achieved with $h=4$, while increasing the truncation threshold yields a noisier estimator.
Figure 4(a) displays the same experiment for data generated with a lower arrival rate of $\lambda=0.3$ (half the utilization rate). This reduces the variance of the estimators and yields CDF estimators with less fluctuations in all cases. As before, the truncation parameter of $h=4$ appears to yield the best fit, although now the difference between the estimated functions is smaller due to the reduced variance. A similar experiment is presented in Figure 4(b), but this time the sampling rate is slower ($\xi=0.2$ instead of $\xi=1$). Again, the optimal choice of truncation is obtained at $h=4$ with slightly lower variance in all cases. The reason for the reduced variance is that the correlation between workload observations is weaker when sampling occurs less frequently. See [31] for a more detailed analysis of the effect of the sampling rate on the variance of estimation errors. Note that a slower sampling rate comes at a cost of a longer sampling duration for the same number of observations. $\diamond$
@@
(a) $\lambda=0.3$, $\xi=1$.
(b) $\lambda=0.6$, $\xi=0.2$.
Figure 4: The estimators are based on a simulation of $n=10^{4}$ workload observations. The plots illustrate the impact of changing the utilization rate and sampling rate. The true CDF $G$ (black solid line) is a log-Normal distribution with $\mu=0.2$ and $\sigma=0.5$. The estimated functions $\widehat{G}_{n}^{(h)}$ are plotted for $h\in\{2,4,6,8\}$..
@@ 3.2 Convergence of the covariance terms
First note that $\mathrm{P}(V_{0}<\infty)=1$ when $V_{0}$ follows the stationary distribution. For an M/G/1 process, it is known that if $\lambda\mathrm{E}[B]<1$ and $\mathrm{E}[B^{k+1}]<\infty$, then the workload process at arrival moments is ergodic with a convergence rate of $n^{k}$ (see [36]). We will show that this result holds for the workload process sampled according to a Poisson process. To this end we first construct a discrete-time version of the workload process at arrival and sampling times. Due to the memoryless property, at any state of the system the time until the next event of an arrival or an observation is an exponential random variable with rate $\lambda+\xi$. Let $(\tau_{k})_{k\geq 0}$ denote an iid sequence of such inter-event times. Let $I_{k}$ denote the indicator of event $k$ being a an arrival event, then $I_{k}$ is a Bernoulli random variable with probability $\lambda/(\lambda+\xi)$, and it is also independent of $\tau_{k}$. Let $\tilde{V}_{k}:=\lim_{t\downarrow 0}V(\sum_{j=1}^{k}\tau_{j}-t)$ denote the workload just before the $k$’th event. The standard Lindley recursion yields
@@
for any bounded function $g$ and initial state $V_{0}\in[0,\infty)$. Let $S(k)$ denote the index of the $k$’th sampling event in the sequence of events driving $(\tilde{V})_{n\geq 1}$, and note that S(k)S(k1)S(k)-S(k-1) follows a geometric distribution with parameter $\xi/(\lambda+\xi)$. By the SLLN, $S(k)/k\>\to_{\rm as}\>\xi/(\lambda+\xi)$ as $k\to\infty$. Clearly, $S(k)\geq k$, and in particular $S(k)\>\to_{\rm as}\>\infty$ as $k\to\infty$. Thus, (38) implies that
@@
Applying L’Hôpital’s rule three times to $\kappa_{1}(s_{1},s_{2};G)$ to establish that the limit as $s_{2}\to 0$ is bounded under the following condition,
@@
$\kappa_{2}(s;G)$ $=\sum_{k=1}^{\infty}\left(\mathrm{E}[\mathrm{e}^{isV_{0}}\mathbf{1}\{V_{k}=0\}]-\mathrm{E}[\mathrm{e}^{isV}]\mathrm{P}(V=0)\right)=\mathrm{E}[\mathrm{e}^{isV}]\mathrm{E}\left[\sum_{k=1}^{\infty}\left(\mathrm{P}(V_{k}=0|V_{0})-\mathrm{P}(V=0)\right)\right]$